QUESTION 5 /4* dx = OA. 4* x(4) In 4 +C In²4 OB. x(4) In4 + 4* In²4 +C ос. x 4 - 4* +C OD. r? (4) 2(In 4) +C о Е. 4 x? (4
Answer #1

Integrating by parts :

\int x4^xdx

=x*integral(4^x)dx - integral(d/dx(x)*integral(4^x)dx)dx

=x(4^x)/ln4 - integral(1*4^x/ln4)dx

=x(4^x)/ln4 - (1/ln4)*integral(4^x)dx

=x(4^x)/ln4 -(1/ln4)*(4^x)/ln4 + C

=x(4^x)/ln4 - 4^x/ln^2(4) + C .

option (A) is correct.